OpenStudy (anonymous):

Assume that y varies inversely with x. If y = 8.8 when x = 2.4, find x when y = 2.1. Round to the nearest tenth.

OpenStudy (anonymous):

2.4/8.8*2.1

OpenStudy (radar):

y=1/kx 8.8=1/2.4k k=.0473 y=i/((.0473)(2.1)) y=10.1

OpenStudy (anonymous):

\[y=\kappa \div \chi\]